OpenStudy (anonymous):

Hi guys I'm trying to get the exact solution of -au''+bu'=x^2, really stuck any help appreciated

zepdrix (zepdrix):

Were you able to find the Complementary Solution at least? I can't help you with the other part :C I just learned about these today, I need to go over them still lol.

OpenStudy (anonymous):

Yeah went through the central differencing and truncation error no problem, but I just seem to be going nowhere with the exact solution

OpenStudy (anonymous):

Okay so what is a function where if you take the derivative and subtract the second derivative, you get \(x^2\)?

OpenStudy (anonymous):

It's probably something like \(\large x^n\) right?

OpenStudy (anonymous):

So let \(u = cx^n\) and let's see if we can't find out what \(c\) and \(n\) are. You following @TKelly

OpenStudy (anonymous):

Yeah sorry just cooking my dinner right now, trying to multitask

OpenStudy (anonymous):

\[bu^\prime = bncx^{n-1}\]\[-au^{\prime \prime} = -a n(n-1)x^{n-2}\]\[x^2=bncx^{n-1}-an(n-1)x^{n-2}\]Something like this...?

OpenStudy (anonymous):

Actually I messed up.

OpenStudy (anonymous):

oh and the boundary conditions are easy u(0)=u(1)=0 I think I've got it from here thanks wio

OpenStudy (anonymous):

Okay let \[u = Ax^2+Bx+C\]\[u^\prime = 2Ax+B\]\[u^{\prime\prime} = 2A\]

OpenStudy (anonymous):

\[x^2 = -a2A + b(2Ax+B)\]Maybe I shouldn't have reused a, but anyway, something like this. @TKelly

OpenStudy (anonymous):

Hmmm this might not work either!! Woops!

OpenStudy (anonymous):

Anyway I hope I helped somewhat.